OpenStudy (anonymous):

simplify 3 over 2 +2 over 3

OpenStudy (anonymous):

.25

OpenStudy (anonymous):

how did u get it?

OpenStudy (anonymous):

3/2 + 2/3 ? =9/6 + 4/6 = 13/6